# OpenRLHF: An Easy-to-use, Scalable and High-performance RLHF Framework

### v5

@@ 4.1 Performance Comparison
To ensure the applicability of the compared methods in current RLHF workflows, we conduct our experimental evaluation under a long-chain-of-thought (CoT) generation scenario. Given computational resource constraints, we focus our comparison on the long CoT RLVR setting, benchmarking OpenRLHF against Verl, currently the state-of-the-art framework for RLHF training. We evaluate the training efficiency of OpenRLHF (v0.8.5) and Verl (v0.4.0) by measuring the average per-step training time (in seconds) across various model sizes (1.5B, 7B, and 14B parameters) and maximum generation lengths (1K, 2K, 4K, and 8K tokens). To ensure the base models can produce sufficiently long contextual outputs for stress testing, we adopt the DeepSeek open-source distilled Qwen series. All models are fine-tuned using the Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) algorithm Yu et al. (2025) under identical hyperparameter settings. Experiments are conducted on 8 NVIDIA H200 140GB GPUs using PyTorch 2.7 Imambi et al. (2021) and the ZeRO Stage 3 optimizer or Fully Sharded Data Parallel (FSDP) Zhao et al. (2023). For each configuration, the local batch size is set to 1 to mitigate the risk of out-of-memory errors, with a maximum input context length of 1024 tokens. The reported values represent the average training time per step, excluding the first 10 steps.
@@
To ensure a fair and controlled evaluation of training efficiency in reinforcement learning fine-tuning, we compare OpenRLHF with the optimized TRL framework on the GSM8K dataset Cobbe et al. (2021) using the GRPO algorithm over a single training epoch. Both systems are configured with identical hyperparameters and run on the same hardware setup to isolate the effect of framework design on performance. The GRPO algorithm is selected due to its relevance in reward-based fine-tuning scenarios, and GSM8K serves as a representative benchmark for arithmetic reasoning tasks. In terms of training efficiency, OpenRLHF demonstrates a substantial advantage, completing one epoch in 1,657 seconds, compared to the 5,189 seconds required by TRL, representing approximately a 3.1× speedup. The result highlights the superior efficiency and maintainability of OpenRLHF’s implementation, which benefits from a streamlined design and targeted system-level optimizations.
@@ 4.2 Usability Comparison
As illustrated in Figure 2, OpenRLHF achieves a compelling balance between implementation conciseness and training performance. Despite being the second most concise framework with only 8,523 lines of code—significantly fewer than TRL (19,071) and Verl (32,325)—OpenRLHF demonstrates a clear performance advantage across standard RLHF benchmarks. This streamlined codebase not only facilitates easier comprehension and modification for developers but also reduces the engineering overhead associated with integration into custom pipelines. In addition to its lightweight design, OpenRLHF offers comprehensive support for various reinforcement learning fine-tuning paradigms, including Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO) Rafailov et al. (2023), Reward Model (RM), and Process Reward Model (PRM) Lightman et al. (2023). This broad functionality, combined with its modular and well-documented architecture, significantly lowers the barrier to entry for both research and production use. Overall, OpenRLHF’s design exemplifies high usability by combining minimal code complexity with extensive functionality and competitive performance.
@@ 5 Limitations
While OpenRLHF demonstrates significant advantages in balancing performance and accessibility, several limitations should be acknowledged. Despite our optimization efforts, OpenRLHF may not match the peak performance of highly specialized industrial frameworks that benefit from dedicated engineering teams and extensive resources. As a community-driven open-source project without dedicated economic support, OpenRLHF faces resource constraints in rapidly integrating cutting-edge features, potentially resulting in delays compared to well-funded commercial frameworks. Currently, the framework focuses primarily on language models and does not support Vision-Language Models or other multimodal architectures, limiting its applicability to multimodal AI alignment research. Additionally, OpenRLHF’s modular design introduces dependencies on external systems such as Ray, vLLM, and DeepSpeed, where updates in these upstream systems may require maintenance work or introduce compatibility issues. Despite these limitations, we believe OpenRLHF’s contributions to accessibility and democratization of RLHF research provide significant value to the community.
@@ 6 Conclusion
We presented OpenRLHF, a simple yet high-performance open-source framework that bridges the gap between performance and usability in RLHF and RLVR training. By integrating Ray’s distributed computing, vLLM’s inference optimization, DeepSpeed ZeRO’s memory efficiency, and ring attention’s sequence parallelism, OpenRLHF delivers four key innovations: Ray-based architecture for streamlined orchestration, 3D parallelism for efficient scaling, accelerated CoT inference addressing the critical bottleneck, and asynchronous dataflow for maximum throughput. The framework’s broad adoption across leading institutions and companies—from academic courses at CMU to production deployments at major tech companies—validates its real-world effectiveness. OpenRLHF’s influence on subsequent frameworks and its role in democratizing RLHF research demonstrate its significant contribution to the field. By open-sourcing this framework, we aim to foster innovation, accelerate research progress, and enable broader participation in aligned AI development.
Appendix A Broad Impact and Adoption
@@
OpenRLHF has been formally integrated into academic curricula, notably being adopted by The CMU Advanced Natural Language Processing Spring 2025 course as a core teaching framework. This integration demonstrates the framework’s educational value and accessibility for students new to RLHF. The framework has also received recognition from major technology communities, including an invitation to present at the PyTorch Expert Exchange 2025, highlighting its technical contributions to the broader deep learning ecosystem.
@@
OpenRLHF’s modular design and extensible architecture have enabled it to serve as a foundation for specialized frameworks. Notable examples include LMM-R1 for multimodal reinforcement learning, MARTI for advanced reasoning tasks, and MM-EUREKA for multi-modal applications. These derivative frameworks demonstrate OpenRLHF’s flexibility in supporting diverse research directions and its role as a platform for innovation in the RLHF ecosystem.
The design principles and technical innovations introduced in OpenRLHF have influenced subsequent framework development in the field. Several prominent frameworks including Verl, Alibaba ROLL, SLIME, and Open-Reasoner-Zero have acknowledged OpenRLHF’s contributions in their documentation and publications, citing its distributed architecture design, Ray-based orchestration approach, and integration strategies as influential to their own development. This acknowledgment reflects OpenRLHF’s role in advancing the state-of-the-art in RLHF system design and establishing best practices for distributed RLHF training.
@@
Appendix B PPO Workflow Design
This section presents OpenRLHF’s comprehensive PPO-based RLHF training workflow, which orchestrates multiple specialized engines to efficiently handle the complex multi-stage training process (as shown in Figure 3). The OpenRLHF PPO workflow consists of four main stages executed iteratively: (1) Rollout Generation, where the current policy generates responses to prompts; (2) Reward Computation, where responses are evaluated using a trained reward model; (3) Advantage Estimation, where advantages and returns are calculated using GAE (Generalized Advantage Estimation); and (4) Policy Optimization, where the policy is updated using PPO loss. This pipeline repeats for multiple iterations until convergence.
The training begins with the Rollout Engine generating responses for a batch of prompts using the current policy $\pi_{\theta}$. A batch of prompts $\{x_{1},x_{2},...,x_{B}\}$ is sampled from the training dataset, and the Rollout Engine, equipped with vLLM for efficient inference, generates responses $\{y_{1},y_{2},...,y_{B}\}$ using the current policy. During generation, the engine records action log-probabilities $\log\pi_{\theta}(y_{i}|x_{i})$ and attention masks. The generated sequences $(x_{i},y_{i})$ along with their metadata are collected for subsequent processing. The Rollout Engine leverages vLLM’s optimizations including continuous batching, KV-cache management, and PagedAttention to maximize throughput during this inference-heavy stage.
@@ Appendix B PPO Workflow Design
Throughout this workflow, OpenRLHF’s Ray-based architecture enables seamless coordination between different engines while maintaining computational efficiency. The Rollout Engine and ZeRO Engine can operate on different GPU clusters optimized for their respective workloads—inference-optimized hardware for rollout generation and training-optimized hardware for policy updates. Ray’s distributed scheduling automatically handles data transfer, synchronization, and fault tolerance across the distributed system. This asynchronous execution model allows the system to overlap computation stages where possible, significantly improving overall training throughput compared to traditional synchronous RLHF implementations.
The entire PPO training loop continues for multiple epochs, with each epoch processing multiple batches of rollout data. Early stopping criteria based on KL divergence thresholds and performance metrics prevent policy collapse and ensure stable training. The modular design allows for easy integration of advanced techniques such as advantage normalization, gradient clipping, and adaptive KL penalty coefficients, making OpenRLHF highly customizable for different research and production scenarios.
Appendix C Full Contributors
  • Ray Integration: Jian Hu, Xibin Wu, Songlin Jiang
  • vLLM Integration: Jian Hu, Xibin Wu, Songlin Jiang, Kaichao You (vLLM Team)
  • Ring Attention: Zilin Zhu (Zhipu), Zhibo Zhou (Vivo), gzpan(GitHub User), Jian Hu
  • Supervised Learning: Haoran Wang, Xianyu
  • DeepSpeed and Transformers Integration: Jian Hu, Xibin Wu, Songlin Jiang, Bin Chen
  • Asynchronous Agentic RL: Haotian Xu, Jian Hu
  • Single Controller Architecture: Jian Hu
  • PPO Implementation: Jian Hu
  • GRPO Implementation: Jason Klein Liu
  • Dynamic Sampling: Jian Hu
  • Dynamic Batching: Hao Chen
  • Documentation: Jian Hu
  • Testing and Bug Reports: Weikai Fang, Wei Shen, Weixun Wang
  • Paper Writing and Presentations: Wei Shen, Jason Klein Liu, Weixun Wang, Yu Cao, Jian Hu

### v6

@@ 4.1 Performance Comparison
To ensure the applicability of the compared methods in current RLHF workflows, we conduct our experimental evaluation under a long-chain-of-thought (CoT) generation scenario. Given computational resource constraints, we focus our comparison on the long CoT RLVR setting, benchmarking OpenRLHF against verl, currently the state-of-the-art framework for RLHF training. We evaluate the training efficiency of OpenRLHF (v0.8.5) and verl (v0.4.0) by measuring the average per-step training time (in seconds) across various model sizes (1.5B, 7B, and 14B parameters) and maximum generation lengths (1K, 2K, 4K, and 8K tokens). To ensure the base models can produce sufficiently long contextual outputs for stress testing, we adopt the DeepSeek open-source distilled Qwen series. All models are fine-tuned using the Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) algorithm [26] under identical hyperparameter settings. Experiments are conducted on 8 NVIDIA H200 140GB GPUs using PyTorch 2.7 [27] and the ZeRO Stage 3 optimizer or Fully Sharded Data Parallel (FSDP) [28]. For each configuration, the local batch size is set to 1 to mitigate the risk of out-of-memory errors, with a maximum input context length of 1024 tokens. Following the discussion in [29], we use $k_{2}$ as the loss function. The reported values represent the average training time per step, excluding the first 10 steps.
@@
To ensure a fair and controlled evaluation of training efficiency in reinforcement learning fine-tuning, we compare OpenRLHF with the optimized TRL framework on the GSM8K dataset [17] using the GRPO algorithm over a single training epoch. Both systems are configured with identical hyperparameters and run on the same hardware setup to isolate the effect of framework design on performance. The GRPO algorithm is selected due to its relevance in reward-based fine-tuning scenarios, and GSM8K serves as a representative benchmark for arithmetic reasoning tasks. In terms of training efficiency, OpenRLHF demonstrates a substantial advantage, completing one epoch in 1,657 seconds, compared to the 5,189 seconds required by TRL, representing approximately a 3.1× speedup. The result highlights the superior efficiency and maintainability of OpenRLHF’s implementation, which benefits from a streamlined design and targeted system-level optimizations.
@@ 4.2 Usability Comparison
As illustrated in Figure 2, OpenRLHF achieves a compelling balance between implementation conciseness and training performance. Despite being the second most concise framework with only 8,523 lines of code—significantly fewer than TRL (19,071) and verl (32,325)—OpenRLHF demonstrates a clear performance advantage across standard RLHF benchmarks. This streamlined codebase not only facilitates easier comprehension and modification for developers but also reduces the engineering overhead associated with integration into custom pipelines. In addition to its lightweight design, OpenRLHF offers comprehensive support for various reinforcement learning fine-tuning paradigms, including Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO) [19], Reward Model (RM), and Process Reward Model (PRM) [30]. This broad functionality, combined with its modular and well-documented architecture, significantly lowers the barrier to entry for both research and production use. Overall, OpenRLHF’s design exemplifies high usability by combining minimal code complexity with extensive functionality and competitive performance.
@@ 5 Limitations
While OpenRLHF demonstrates significant advantages in balancing performance and accessibility, several limitations should be acknowledged. Despite our optimization efforts, OpenRLHF may not match the peak performance of highly specialized industrial frameworks that benefit from dedicated engineering teams and extensive resources. As a community-driven, open-source project without dedicated economic support, OpenRLHF faces resource constraints in rapidly integrating cutting-edge features, which may result in delays compared to well-funded commercial frameworks. Currently, the framework primarily focuses on language models and does not support Vision-Language Models or other multimodal architectures, thereby limiting its applicability to multimodal AI alignment research. Additionally, OpenRLHF’s modular design introduces dependencies on external systems such as Ray, vLLM, and DeepSpeed, where updates in these upstream systems may require maintenance work or introduce compatibility issues. Despite these limitations, we believe OpenRLHF’s contributions to accessibility and democratization of RLHF research provide significant value to the community.
@@ 6 Conclusion
We presented OpenRLHF, a simple yet high-performance open-source framework that bridges the gap between performance and usability in RLHF and RLVR training. By integrating Ray’s distributed computing, vLLM’s inference optimization, DeepSpeed ZeRO’s memory efficiency, and ring attention’s sequence parallelism, OpenRLHF delivers four key innovations: a Ray-based architecture for streamlined orchestration, 3D parallelism for efficient scaling, accelerated CoT inference that addresses the critical bottleneck, and asynchronous dataflow for maximum throughput. The framework’s broad adoption across leading institutions and companies—from academic courses at CMU to production deployments at major tech companies—validates its real-world effectiveness. OpenRLHF’s influence on subsequent frameworks and its role in democratizing RLHF research demonstrate its significant contribution to the field. By open-sourcing this framework, we aim to accelerate research progress and enable broader participation in the development of aligned AI.
Appendix A Full Contributors
Ray Integration
Jian Hu, Xibin Wu, Songlin Jiang
vLLM Integration
Jian Hu, Xibin Wu, Songlin Jiang, Kaichao You (vLLM Team)
Ring Attention
Zilin Zhu (Zhipu), Zhibo Zhou (Vivo), gzpan(GitHub User), Jian Hu
Supervised Learning:
Haoran Wang and Xianyu
DeepSpeed Integration
Jian Hu, Xibin Wu, Songlin Jiang, Bin Chen, Yiming Liu
Asynchronous Agentic RL
Haotian Xu and Jian Hu
Single Controller Architecture
Jian Hu
PPO Implementation
Jian Hu
GRPO Implementation
Jason Klein Liu
KL Control Mechanism
Yiming Liu and Jason Klein Liu
RLVR Experiment Design
Jason Klein Liu
RLHF Experiment Design
Jian Hu
Dynamic Sampling
Jian Hu
Dynamic Batching
Hao Chen
Documentation
Jian Hu
Testing and Bug Reports
Weikai Fang, Wei Shen, Weixun Wang
Paper Writing and Presentations
Wei Shen, Jason Klein Liu, Weixun Wang, Yu Cao, Jian Hu
Appendix B Broad Impact and Adoption
@@
OpenRLHF has been formally integrated into academic curricula, notably being adopted by the CMU Advanced Natural Language Processing course in Spring 2025 as a core teaching framework. This integration demonstrates the framework’s educational value and accessibility for students new to RLHF. The framework has also garnered recognition from major technology communities, including an invitation to present at the PyTorch Expert Exchange 2025, which highlights its technical contributions to the broader deep learning ecosystem.
@@
OpenRLHF’s modular design and extensible architecture have enabled it to serve as a foundation for specialized frameworks. Notable examples include LMM-R1 for multimodal reinforcement learning, MARTI for advanced reasoning tasks, and MM-EUREKA for multimodal applications. These derivative frameworks demonstrate OpenRLHF’s flexibility in supporting diverse research directions and its role as a platform for innovation in the RLHF ecosystem.
The design principles and technical innovations introduced in OpenRLHF have influenced subsequent framework development in the field. Several prominent frameworks, including verl, Alibaba ROLL, SLIME, and Open-Reasoner-Zero, have acknowledged OpenRLHF’s contributions in their documentation and publications, citing its distributed architecture design, Ray-based orchestration approach, and integration strategies as influential to their own development. This acknowledgment reflects OpenRLHF’s role in advancing the state-of-the-art in RLHF system design and establishing best practices for distributed RLHF training.
@@
Appendix C PPO Workflow Design
This section presents OpenRLHF’s comprehensive PPO-based RLHF training workflow, which orchestrates multiple specialized engines to efficiently handle the complex multi-stage training process (as shown in Figure 3). The OpenRLHF PPO workflow consists of four main stages executed iteratively: (1) Rollout Generation, where the current policy generates responses to prompts; (2) Reward Computation, where responses are evaluated using a trained reward model; (3) Advantage Estimation, where advantages and returns are calculated using GAE (Generalized Advantage Estimation); and (4) Policy Optimization, where the policy is updated using PPO loss. This pipeline repeats for multiple iterations until convergence is achieved.
The training begins with the Rollout Engine generating responses for a batch of prompts using the current policy $\pi_{\theta}$. A batch of prompts $\{x_{1},x_{2},...,x_{B}\}$ is sampled from the training dataset, and the Rollout Engine, equipped with vLLM for efficient inference, generates responses $\{y_{1},y_{2},...,y_{B}\}$ using the current policy. During generation, the engine records action log-probabilities $\log\pi_{\theta}(y_{i}|x_{i})$ and attention masks. The generated sequences $(x_{i},y_{i})$ along with their metadata are collected for subsequent processing. The Rollout Engine leverages vLLM’s optimizations, including continuous batching, KV-cache management, and PagedAttention, to maximize throughput during this inference-heavy stage.
@@ Appendix C PPO Workflow Design
Throughout this workflow, OpenRLHF’s Ray-based architecture enables seamless coordination between different engines while maintaining computational efficiency. The Rollout Engine and ZeRO Engine can operate on different GPU clusters, each optimized for their respective workloads: inference-optimized hardware for rollout generation and training-optimized hardware for policy updates. Ray’s distributed scheduling automatically handles data transfer, synchronization, and fault tolerance across the distributed system. This asynchronous execution model enables the system to overlap computation stages whenever possible, thereby significantly improving overall training throughput compared to traditional synchronous RLHF implementations.
The entire PPO training loop continues for multiple epochs, with each epoch processing multiple batches of rollout data. Early stopping criteria based on KL divergence thresholds and performance metrics prevent policy collapse and ensure stable training. The modular design enables the easy integration of advanced techniques, such as advantage normalization, gradient clipping, and adaptive KL penalty coefficients, making OpenRLHF highly customizable for various research and production scenarios.